The gain of an amplifier is found to be 𝐺 = 20 ln(π‘‰π‘œπ‘’π‘‘). The tasks are to find equations for: a) Draw a graph of Gain against Vout between π‘‰π‘œπ‘’π‘‘ = 1 and π‘‰π‘œπ‘’π‘‘ = 10 b) Determine the gradient of the graph at π‘‰π‘œπ‘’π‘‘ = 2 and π‘‰π‘œπ‘’π‘‘ = 5 c) Find the derivative 𝑑𝐺/ 𝑑𝑉𝑂𝑒𝑑 and calculate its value at π‘‰π‘œπ‘’π‘‘ = 2 and π‘‰π‘œπ‘’π‘‘ = 5 d) Compare your answers for part b) and part c) e) Find the second derivative 𝑑2𝐺/ 𝑑𝑉𝑂𝑒𝑑 2Β 


1
Expert's answer
2022-01-04T10:55:34-0500

Solution:

Given 𝐺 = 20 ln(π‘‰π‘œπ‘’π‘‘)

From π‘‰π‘œπ‘’π‘‘ = 1 and π‘‰π‘œπ‘’π‘‘ = 10.

Treat π‘‰π‘œπ‘’π‘‘=x and G=y

(a):

then, the graph is:



(b):

Gradient At π‘‰π‘œπ‘’π‘‘ =2= slope = "\dfrac yx= \dfrac{20 \ln(2)}{2}=6.931"

Gradient At π‘‰π‘œπ‘’π‘‘ =5= slope = "\dfrac yx= \dfrac{20 \ln(5)}{5}=6.437"

(c):

"\dfrac{d𝐺}{d 𝑉_{𝑜𝑢𝑡}} = 20\times \dfrac{1}{𝑉_{𝑜𝑢𝑡}} =\dfrac{20}{𝑉_{𝑜𝑢𝑡}}"

At π‘‰π‘œπ‘’π‘‘ =2, "\dfrac{d𝐺}{d 𝑉_{𝑜𝑢𝑡}} =\dfrac{20}{2}=10"

At π‘‰π‘œπ‘’π‘‘ =5, "\dfrac{d𝐺}{d 𝑉_{𝑜𝑢𝑡}} =\dfrac{20}{5}=4"

(d):

There is huge difference in answers in part b and c.

(e):

"\dfrac{d𝐺}{d 𝑉_{𝑜𝑢𝑡}} =\dfrac{20}{𝑉_{𝑜𝑢𝑡}}
\\ \Rightarrow \dfrac{d^2G}{d 𝑉_{𝑜𝑢𝑡}^2} =\dfrac{-20}{𝑉_{𝑜𝑢𝑡}^2}"